OpenStudy (anonymous):

what are the solutions to the following systems of equations? y=x^2-8x+2 y=-x^2-10x+14

OpenStudy (anonymous):

Set them equal to each other and then Simplify

OpenStudy (anonymous):

\[x^2-8x+2 = x^2-10x+14\] solve for x

OpenStudy (anonymous):

there is a minus after the equal sign http://www.wolframalpha.com/input/?i=solve+x^2+-+8+x+%2B+2+%3D+-x^2+-+10+x+%2B+14

OpenStudy (anonymous):

oops, I just copied and pasted. must have missed that

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=x^2-8x+2 \qquad {\color{red}{ y}}=-x^2-10x+14 \\ \quad \\ \implies x^2-8x+2=-x^2-10x+14\implies 2x^2+2x-12=0 \\ \quad \\ 2(x^2+x-6)=0\) solve for "x"